OpenStudy (anonymous):

Find the range of values for x for which the equation x^2 - 2k(x+1) +k^2 = 6 has real roots. Find the roots in terms of k.

OpenStudy (zzr0ck3r):

x^2 -2kx-2k+k^2-6=0 so sqrt(b^2-4ac) must be real sqrt((2k)^2-4*(-2k+k^2) so ((2k)^2-4*(-2k+k^2)>=0 4k^2+8k-4k-4k^2 4k>=0 so k>=0

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

wait a min

OpenStudy (anonymous):

x^2 -2kx-2k+k^2-6=0 so sqrt(b^2-4ac) must be real sqrt((2k)^2-4*(-2k+k^2) <-- where did the -6 go lol

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

good call, one sec

OpenStudy (zzr0ck3r):

((2k)^2-4*(-2k+k^2-6)>=0 4k^2+8k-4k^2+24>=0 8k+24>=0 8k>=-24 k>=-3 OK im glad you saw that because my answer was not working when I try it.

OpenStudy (zzr0ck3r):

and this one does when I check it:)

OpenStudy (anonymous):

thanks!

OpenStudy (zzr0ck3r):

find roots x^2 -2kx-2k+k^2-6=0 (x-k)^2 = 2k-k^2+6+k^2 x = +_sqrt(2k-k^2+6+k^2)+k

OpenStudy (zzr0ck3r):

you see what I did?

OpenStudy (zzr0ck3r):

I completed the square.

OpenStudy (anonymous):

yup alright

OpenStudy (zzr0ck3r):

should that say find the values of K for wich the equation has a solution?

OpenStudy (zzr0ck3r):

which*

OpenStudy (zzr0ck3r):

@BelleFlower ?

OpenStudy (anonymous):

nope i copied the question correctly..

OpenStudy (zzr0ck3r):

doesn't make sense.

OpenStudy (zzr0ck3r):

that like what values of x make x^2+2x+1=0 have real solutions real solutions don't depend on the variable they depend on the constants. and K is the constand

OpenStudy (zzr0ck3r):

constant*

OpenStudy (zzr0ck3r):

I think it meant K, because what we did makes sense...

OpenStudy (zzr0ck3r):

lol bye